# Model-Free Assessment of Simulator Fidelity via Quantile Curves

### v2

@@ 6 Discussion
Despite its broad applicability, the method we introduce leaves several natural extensions. First, our proof relies on DKW-type concentration – often conservative for small $m$ – and a grid-uniform step that further loosens constants; tightening these bounds is an immediate priority. Second, many applications involve temporally dependent, dynamic simulation processes; extending our static framework to dynamic settings would broaden applicability. Third, our analysis assumes i.i.d. scenarios, whereas covariate shift or endogenous sampling may invalidate marginal guarantees; addressing such distribution shifts is an important avenue for future work.
@@ Appendix A Technical Lemmas
$\mathbb{P}_{\theta}(\bar{T}_{n}\geq m)\leq\exp\!\big(-n\psi_{\theta}^{*}(m)\big),$
$\implies$ $\psi_{\theta}^{*}(m):=\sup_{\lambda\in\mathbb{R}}\{\lambda m-\psi_{\theta}(\lambda)\},$
@@
$\bar{F}_{m}(t)$ $:=\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}\{\Delta_{j}\leq t\},$
$\bar{V}_{m}(u)$ $:=\inf\{t:\bar{F}_{m}(t)\geq u\}.$
@@
$\mathbb{P}_{\psi\sim\Psi}(L(p(\psi),\hat{q}(\psi))$ $\leq\bar{V}_{m}(1-\frac{\bar{\alpha}}{2})|\mathcal{D})$
$\geq 1-\frac{\bar{\alpha}}{2}-\sqrt{\frac{\log(2/\delta)}{2m}}$ (B.6)
@@
$\mathbb{P}\Big(|T_{\alpha}|\leq\gamma s-t|\mathcal{G}\Big)$ $\leq\mathbb{P}\Big(|T_{\alpha}|\leq\mathbb{E}[T_{\alpha}]-t|\mathcal{G}\Big)$
$\leq\exp(-\frac{2t^{2}}{s}),$
@@
$y^{*}$ $=\frac{c/\sqrt{m}+\sqrt{c^{2}/m+4\gamma\alpha+4\gamma/m}}{2\gamma},$
$\alpha_{\mathrm{eff}}(\alpha,c,m)$ $=(y^{*})^{2}.$
@@
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}\leq\bar{V}_{m}(1-\alpha_{\mathrm{eff}})\Big|\mathcal{D}\Big)$
$\geq 1-\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)-\varepsilon_{m}(\delta).$
@@
$\alpha_{\mathrm{eff}}(\bar{\alpha},$ $c,m)=\frac{\big(\tfrac{c}{\sqrt{m}}+\sqrt{\tfrac{c^{2}}{m}+4\gamma\bar{\alpha}+\tfrac{4\gamma}{m}}\big)^{2}}{4\gamma^{2}}=$
$\frac{\bar{\alpha}}{\gamma}+\frac{c}{\gamma\sqrt{m}}\sqrt{4\gamma\bar{\alpha}+\tfrac{c^{2}+4\gamma}{m}}+\frac{c^{2}}{2\gamma^{2}m}+\frac{1}{\gamma m}.$
@@
$\mathbb{P}_{\psi\sim\Psi}$ $\Big(\Delta(\psi)\leq\widehat{V}_{m}(1-\alpha)\Big|\mathcal{D}\Big)$
$\geq 1$ $-\frac{\alpha_{+}}{\gamma}-\frac{c}{\gamma\sqrt{m}}\sqrt{4\gamma\alpha_{+}+\tfrac{c^{2}+4\gamma}{m}}$
$-\frac{c^{2}}{2\gamma^{2}m}-\frac{1}{\gamma m}-\varepsilon_{m}(\delta).$
@@ E.2 OpinionQA
Our dataset is OpinionQA Santurkar et al. (2023b), built from the Pew Research’s American Trends Panel, which consists of the US population’s responses to survey questions spanning topics such as racial equity, security, and technology. We adopt the preprocessed version curated by Huang et al. (2025), which includes 385 distinct questions and 1,476,868 responses from at least 32,864 people. Each question has 5 choices, corresponding to the order sentiments, which is a multinomial setting. We can construct confidence sets $\mathcal{C}_{j}$ for multinomial vectors by adopting Example 3.1 with $d=5$. OpinionQA also provides individual-level covariates such as gender, age, socioeconomic status, religious affiliation, and marital status, and more, which are used to construct synthetic profiles. Under the same problem formulation, the authors of Huang et al. (2025) compute $\{\hat{p}_{j},\hat{q}_{j}\}_{j=1}^{385}$ for seven LLMs: GPT-3.5-Turbo (gpt-3.5-turbo), GPT-4o (gpt-4o), and GPT-4o-mini (gpt-4o-mini); Claude 3.5 Haiku (claude-3-5-haiku-20241022); Llama 3.3 70B (Llama-3.3-70B-Instruct-Turbo); Mistral 7B (Mistral-7B-Instruct-v0.3); DeepSeek-V3 (DeepSeek-V3), and constructed a baseline random simulator that selects uniformly among the available answer choices. A more detailed exploration into the OpinionQA dataset and the simulation procedure can be found in Huang et al. (2025). We also provide the same procedure onto a Bernoulli setting using the EEDI dataset, details can be found in Appendix E.

### v3

@@ 6 Discussion
Despite its broad applicability, the method we introduce leaves several natural extensions. First, our framework targets scenario-level population discrepancy: for each scenario, the object of interest is a marginal population-level response distribution or derived parameter. As a result, the current formulation does not model respondent-level joint dependence across multiple questions. Second, our guarantees are developed for static per-scenario outputs under an i.i.d. scenario assumption, and do not yet cover fully dynamic simulators with temporal dependence, endogenous sampling, or distribution shift. Third, on the computational side, the main bottleneck is simulator-side generation: once real and simulated batches are available, our methodology computationally is lightweight, but expensive simulators may make sample generation itself the limiting factor. Finally, because the method relies on generic concentration inequalities and worst-case pseudo-discrepancies, the resulting envelope can be conservative when the number of scenarios $m$ or the per-scenario sample sizes are small. In particular, our proof relies on DKW-type concentration—which can be loose for small mm—together with a grid-uniform step that further weakens constants, so tightening these bounds is an immediate priority.
@@ Appendix A Technical Lemmas
$\mathbb{P}_{\theta}(\bar{T}_{n}\geq m)\leq\exp\big(-n\psi_{\theta}^{*}(m)\big),\implies\psi_{\theta}^{*}(m):=\sup_{\lambda\in\mathbb{R}}\{\lambda m-\psi_{\theta}(\lambda)\},$
@@
$\bar{F}_{m}(t):=\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}\{\Delta_{j}\leq t\},\quad\bar{V}_{m}(u):=\inf\{t:\bar{F}_{m}(t)\geq u\}.$
@@
$\mathbb{P}_{\psi\sim\Psi}(L(p(\psi),\hat{q}(\psi))\leq\bar{V}_{m}(1-\frac{\bar{\alpha}}{2})|\mathcal{D})\geq 1-\frac{\bar{\alpha}}{2}-\sqrt{\frac{\log(2/\delta)}{2m}}$ (B.6)
@@
$\mathbb{P}\Big(|T_{\alpha}|\leq\gamma s-t|\mathcal{G}\Big)\leq\mathbb{P}\Big(|T_{\alpha}|\leq\mathbb{E}[T_{\alpha}]-t|\mathcal{G}\Big)\leq\exp(-\frac{2t^{2}}{s}),$
@@
$y^{*}=\frac{c/\sqrt{m}+\sqrt{c^{2}/m+4\gamma\alpha+4\gamma/m}}{2\gamma},\quad\alpha_{\mathrm{eff}}(\alpha,c,m)=(y^{*})^{2}.$
@@
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}\leq\bar{V}_{m}(1-\alpha_{\mathrm{eff}})\Big|\mathcal{D}\Big)\geq 1-\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)-\varepsilon_{m}(\delta).$
@@
$\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)$ $=\frac{\big(\tfrac{c}{\sqrt{m}}+\sqrt{\tfrac{c^{2}}{m}+4\gamma\bar{\alpha}+\tfrac{4\gamma}{m}}\big)^{2}}{4\gamma^{2}}$
$=\frac{\bar{\alpha}}{\gamma}+\frac{c}{\gamma\sqrt{m}}\sqrt{4\gamma\bar{\alpha}+\tfrac{c^{2}+4\gamma}{m}}+\frac{c^{2}}{2\gamma^{2}m}+\frac{1}{\gamma m}.$
@@
$\mathbb{P}_{\psi\sim\Psi}$ $\Big(\Delta(\psi)\leq\widehat{V}_{m}(1-\alpha)\Big|\mathcal{D}\Big)\geq 1-\frac{\alpha_{+}}{\gamma}-\frac{c}{\gamma\sqrt{m}}\sqrt{4\gamma\alpha_{+}+\tfrac{c^{2}+4\gamma}{m}}-\frac{c^{2}}{2\gamma^{2}m}-\frac{1}{\gamma m}-\varepsilon_{m}(\delta).$
@@ Appendix D Simulation System Examples
Building control: room-temperature response (stochastic emulator; short-horizon thermal dynamics).
  • Outcome space: $\mathcal{X}=\mathbb{R}$, where the outcome is a short-horizon room-temperature response.
  • Scenarios: $\psi$ denotes a coarse operational context across which simulator fidelity is compared, for example building or unit identity together with hour-of-day, day-type, and an ambient-temperature regime.
  • Profiles: $\mathcal{Z}$ collects the within-scenario thermal state and disturbance variables that vary even after $\psi$ is fixed, such as current and lagged room temperature, solar irradiance, current heating setpoint, and other occupant- or disturbance-related variation.
  • Laws: $Q^{\mathrm{gt}}(\cdot\mid z,\psi)$ denotes the real conditional law of future room temperature given the current profile $z$ and operational context $\psi$. $Q^{\mathrm{sim}}(\cdot\mid a(z),\psi,r)$ denotes the emulator output under mirrored pre-decision inputs $a(z)$, context $\psi$, and simulator randomness $r$.
  • Parameters: $\Theta=\mathbb{R}$ if the target is the scenario-level mean future room temperature, or $\Theta=\mathbb{R}^{2}$ if one summarizes the law by mean and variance; alternatively, one may target a small collection of temperature quantiles.
  • Discrepancy: $L$ may be the absolute difference of scenario-level means, a Wasserstein distance on the scalar room-temperature distribution, or a Gaussian Kullback–Leibler divergence when a parametric approximation is adopted.
  • Sampling: $n_{j}$ real windows observed in scenario $j$, each contributing one realized future room-temperature outcome; $k$ emulator draws generated under the same scenario.
@@ E.2 OpinionQA
Our dataset is OpinionQA Santurkar et al. (2023b), built from the Pew Research’s American Trends Panel, which consists of the US population’s responses to survey questions spanning topics such as racial equity, security, and technology. We adopt the preprocessed version curated by Huang et al. (2025), which includes 385 distinct questions and 1,476,868 responses from at least 32,864 people. Each question has 5 choices, corresponding to the order sentiments, which is a multinomial setting. We can construct confidence sets $\mathcal{C}_{j}$ for multinomial vectors by adopting Example 3.1 with $d=5$. OpinionQA also provides individual-level covariates such as gender, age, socioeconomic status, religious affiliation, and marital status, and more, which are used to construct synthetic profiles. Under the same problem formulation, the authors of Huang et al. (2025) compute $\{\hat{p}_{j},\hat{q}_{j}\}_{j=1}^{385}$ for eight LLMs: GPT-3.5-Turbo (gpt-3.5-turbo), GPT-4o (gpt-4o), and GPT-4o-mini (gpt-4o-mini); Claude 3.5 Haiku (claude-3-5-haiku-20241022); Llama 3.3 70B (Llama-3.3-70B-Instruct-Turbo); Llama 3.1 8B (Llama-3-8B-InstructTurbo) ; Mistral 7B (Mistral-7B-Instruct-v0.3); DeepSeek-V3 (DeepSeek-V3), and constructed a baseline random simulator that selects uniformly among the available answer choices. A more detailed exploration into the OpinionQA dataset and the simulation procedure can be found in Huang et al. (2025). We also provide the same procedure onto a Bernoulli setting using the EEDI dataset, details can be found in Appendix E.
@@
E.3 UMAR Dataset
We consider the UMAR unit of NEST (Next Evolution in Sustainable Building Technologies), a modular living-lab platform for building research operated by Empa Richner et al. (2018), which is released by Heer et al. (2024). The published dataset contains four years of one-minute measurements, from July 1, 2019 to June 30, 2023, and includes room temperatures, HVAC setpoints, valve and window states, ambient temperature, solar irradiation, and building-energy variables. In our reproduction, the UMAR portion is stored as four yearly wide-format CSV files together with a metadata table for the UMAR measurement channels.
Following the fixed experimental design used in our UMAR reproduction, we focus on room temp_275 and aggregate the one-minute measurements to 30-minute means. We then construct a prediction table using the current room temperature, two lagged temperatures, hour of day, weekend indicator, month, ambient temperature, and, when available, room setpoint and solar irradiation as covariates. To avoid clear sensor artifacts and outliers, we remove rows for which room_temp_t, lag1, lag2, or the future response exceeds $30^{\circ}$C. This yields 66,922 cleaned 30-minute observations.
Unlike EEDI and OpinionQA, UMAR is not a multinomial-response setting. Here the unit of analysis is an operational scenario rather than a question. We define the scenario label as
$\psi=(h,w,t_{q}),$
where $h\in\{0,\dots,23\}$ is the hour of day, $w\in\{0,1\}$ indicates whether the timestamp falls on a weekend, and $t_{q}\in\{0,1,2,3\}$ is the quartile of the ambient temperature. This gives $m=192$ scenarios in total. For each scenario $\psi_{j}$, the real-side parameter $p_{j}$ is the variance of the next-step 30-minute room temperature among held-out UMAR observations in that scenario, while the simulator-side parameter $q_{j}$ is the corresponding variance induced by emulator draws. Since this is a continuous-outcome problem, we construct the confidence set $\mathcal{C}_{j}$ for $p_{j}$ using a bootstrap confidence interval for the scenario-level variance, with confidence level $\gamma_{j}=1-n_{j}^{-1/3}$, where $n_{j}$ is the number of real observations in scenario $j$.
We compare three emulator families using five-fold cross-fitting: linear regression model, decision tree, and multilayer perceptron. For each emulator and each scenario, we generate a balanced simulator budget of $k=200$ draws. We then apply our quantile-fidelity procedure to the resulting scenario-level variance discrepancies.
Figure 15: Adjusted upper quantile curves for scenario-level variance discrepancy across UMAR emulator families. Lower curves indicate closer agreement between emulator-side and real-side temperature variability.
Figure 15 reports the adjusted upper quantile curves for the UMAR experiment. Lower and flatter curves indicate that the emulator more faithfully reproduces the scenario-level variability observed in the real building data. In this reproduction, the linear regression model attains the lowest overall upper quantile curve, with the multilayer perceptron close behind, while the decision-tree baseline exhibits substantially larger upper-tail discrepancies. This illustrates that predictive accuracy alone is not the main object of interest here: the fidelity target is the distribution of scenario-level variability, not only pointwise temperature prediction error.